OpenStudy (anonymous):

A triangle has a height of 2x and a base of 7-4x. 1. What is the maximum area of the triangle? 2. What value of x gives the maximum area?

OpenStudy (campbell_st):

well find the area \[A = \frac{1}{2} \times 2x \times (7 - x)\]. What do you get?

OpenStudy (campbell_st):

to find the maximum area. find the 1st derivative and then let it equal zero and solve for x. This will be the value of x that creates the maximum area.

OpenStudy (campbell_st):

the Alternate method, if you don't know calculus is graph the area curve. the highest point of the parabola is the max area.

OpenStudy (campbell_st):

you can graph the curve using https://www.desmos.com/calculator

OpenStudy (campbell_st):

another alternative is, once you have the Area equation. Find the line of symmetry, for a curve \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\] then substitute thus value into the Area equation to find the max area

OpenStudy (anonymous):

if i simplify A=(1/2)(2x)(7−x) would i get 7x-4x^2 ? from that would i just complete the square to find the max?

OpenStudy (campbell_st):

no you can differentiate \[\frac{dA}{dx} = 7 - 8x\] solve for x x = 7/8 so the max area occurs when x = 7/8 units.... just substitute to find the area.